# Gauge fields induced by curved spacetime

### v2

@@ Abstract
Equation 4 is, in essence, a discrete tight-binding Hamiltonian with lattice-dependent hopping amplitudes. As such, there are several physical systems that simulate this model, e.g., arrays of lattices of atoms deposited on a surface[39, 40], arrays of quantum dots[41], cold atoms in optical lattices[42, 43, 44, 45, 46, 47], photonic crystals[16], superconducting quantum circuits[48, 49, 50], topologically nontrivial stripes[51], and exciton-polariton condensates in artificial lattices[52]. Controlling the hopping amplitudes corresponds to controlling the distances and overlap amplitudes between contiguous localized states on the lattice.
@@
Figure 2(a) show the energy spectra of the Hamiltonian $\mathcal{H}_{\text{CS}}$ (Hofstadter butterfly) describing massless Dirac fermions in curved spacetime on a lattice as a function of the wavelength $\omega$ of the periodic spacetime metric. Figures 2(b) to 2(d) show the energy spectrum and edge modes calculated with open boundary conditions for $p/q=1/3$ and the corresponding periodic spacetime metric in Eq. 12. Figures 2(e) to 2(g) show the energy spectrum and some of the bulk modes calculated with periodic boundary conditions and the corresponding quasiperiodic spacetime metric in Eq. 12 for a wavelength incommensurate with the lattice, specifically $\omega=2\pi(\Phi-1)$, where $\Phi$ is the golden ratio. Note that the dispersion of the energy bands in the synthetic dimension $\phi$ is flattened, which corresponds to energy modes delocalized on the entire lattice. Note also that the spacetime metric is not periodic anymore and shows large amplitude oscillations.

### v3

@@ V Discussion
Equation 4 is, in essence, a discrete tight-binding Hamiltonian with lattice-dependent hopping amplitudes. As such, there are several physical systems that simulate this model, e.g., arrays of lattices of atoms deposited on a surface [42, 43], arrays of quantum dots [44], cold atoms in optical lattices [45, 46, 47, 48, 49, 50], photonic crystals [19], superconducting quantum circuits [51, 52, 53], topologically nontrivial stripes [54], and exciton-polariton condensates in artificial lattices [55]. Controlling the hopping amplitudes corresponds to controlling the distances and overlap amplitudes between contiguous localized states on the lattice. Lattice models of curved spacetime provide ways to simulate gravitational effects in condensed matter, such as the Hawking/Unruh effect [56, 57, 58, 59, 60] or AdS/CFT [61, 62] or to use geometric language to describe condensed matter phenomena [63, 16].
@@
Figure 2(a) shows the energy spectra of the Hamiltonian $\mathcal{H}_{\text{CS}}$ (Hofstadter butterfly) describing massless Dirac fermions in curved spacetime on a lattice as a function of the wavelength $\omega$ of the periodic spacetime metric. Figures 2(b) to 2(d) show the energy spectrum and edge modes calculated with open boundary conditions for $p/q=1/3$ and the corresponding periodic spacetime metric in Eq. 14. Figures 2(e) to 2(g) show the energy spectrum and some of the bulk modes calculated with periodic boundary conditions and the corresponding quasiperiodic spacetime metric in Eq. 14 for a wavelength incommensurate with the lattice, specifically $\omega=2\pi(\Phi-1)$, where $\Phi$ is the golden ratio. Note that the dispersion of the energy bands in the synthetic dimension $\phi$ is flattened, corresponding to energy modes delocalized on the entire lattice. Note also that the spacetime metric is not periodic anymore and shows large amplitude oscillations. Figures 2(h) to 2(j) show the energy spectrum and some of the bulk modes with the corresponding spacetime metric for a small wavelength $\omega=2\pi/N$, with $N$ the number of lattice sites. Notice that this metric becomes linear at the center $x=N/2$ of the system.
Figure 2: Energy spectra of the Hamiltonian $\mathcal{H}_{\text{CS}}$ of a massless Dirac fermion in a periodic spacetime metric on a lattice of $N=210$ sites. (a) Energy spectra as a function of $p/q$ with periodic boundary conditions and $\phi=0$. (b) Energy spectra for $p/q=1/3$ as a function of the phase $\phi$ with open boundary conditions. (c) The metric $\alpha_{n}$ for $p/q=1/3$ and $\phi=0$ compared with $\cos{\left(\frac{1}{2}(\omega n+\phi)\right)}$ ) (dashed) as a function of the lattice site. (d) Wavefunctions of the two edge modes in the first band gap (from below) for $p/q=1/3$ and $\phi=0$. (e) Energy spectra for $\omega\approx 2\pi(\Phi-1)$ (with $\Phi$ the golden ratio), as a function of the phase $\phi$ with periodic boundary conditions on a lattice of $N=233$ sites. (f) The metric $\alpha_{n}$ for $\omega\approx 2\pi(\Phi-1)$ and $\phi=0$ as a function of the lattice site. (g) Wavefunctions of two bulk modes for $\omega\approx 2\pi(\Phi-1)$ and $\phi=0$. (h) Energy spectra for small $\omega\approx 0$, as a function of the phase $\phi$ with periodic boundary conditions on a lattice of $N=233$ sites. The spectra show several discrete levels at a short distance, which corresponds to continuum of energy levels in the limit $\omega\to 0$. (i) The metric $\alpha_{n}$ for $\omega\approx 0$ and $\phi=0$ as a function of the lattice site. Notice that this metric becomes linear at the center $x=N/2$. (j) Wavefunctions of two bulk modes for $\omega\approx 0$ and $\phi=0$.
Note that the metric in Eq. 7 yields the familar Rindler metric if $\alpha(x)=rx$. The Rindler metric is the metric of a reference frame in uniform acceleration in flat spacetime, and has a singularity at $x=0$, where the metric tensor has zero determinant and approximates the Schwarzschild metric near a black hole horizon. This metric is approximately realized by Eq. 14 for small frequencies $\omega$ in finite patches of the spacetime in the case where the metric does not exhibit divergencies [see, e.g., Fig. 2(i)]. This setup allows to simulate a black hole event horizon and in particular the Unruh effect, i.e., the emergence of a thermal distribution of excited states at zero-temperature induced by the presence of the event horizon [64, 65, 66, 67, 68], with the resulting a nonzero effective temperature analogous of Hawking or Unruh radiation [65, 67]. In this setup the Unruh radiation is produceed by a sudden quench corresponding to an abrupt change in the hopping term (i.e., the spacetime curvature). Note that, these phenomena are not experimentally accessible in high-energy experiments, since the Universe we live in is locally flat. Moreover, by allowing a time-dependence of the spacetime curvature, the Hamiltonian $\mathcal{H}_{\text{CS}}$ allows to simulate dilating and contracting universes, which is again not direclty observable in our universe timeline.
VI Conclusions
@@
On a more fundamental level, the triality can lead to novel ways to look at quantum gravity theories. Notice that the spacetime curvature considered here is classical (not quantized) while the fermion fields are quantized. One may consider quantized gravity models such as those studied in the context of quantum gravitational lensing in Ref. [69, 70] by promoting the curvature component to operators $\alpha(x)\to\alpha(\hat{x})$. Another approach is to consider entangled states induced by the superpositions of two curvatures corresponding, e.g., to the Newtonian potential in a Schrödinger-Newton gravity model. This leads to a condensed matter analog of quantum gravity, which is in turn equivalent to a system with flat curvature and a finite gauge field, due to the triality. Hence, the triality opens the door to studying quantized gravity models via studying the better-understood gauge fields in QED. This may hint at novel directions in pursuing a unified theory of quantum mechanics and general relativity.
@@ Acknowledgements.
Appendix A Analytical demonstration that the metric is periodic for commensurate wavelengths
Here, I demonstrate that
${\alpha_{n}}=C^{(-1)^{n}}\prod_{m=1}^{n-1}\left[\cos^{2}\left(\frac{1}{2}(\omega m+\phi)\right)\right]^{(-1)^{m+n+1}},$ (14)
is a periodic function of the lattice index $n$ with period $2q$ when $\omega=2\pi p/q$ with $p,q$ coprimes and $q$ odd. Assume $C=1$, and notice that for $n>2q$ one has
${\alpha_{n+2q}}=$ $\prod_{m=1}^{2q}\left[\cos^{2}\left(\frac{pm\pi}{q}+\frac{\phi}{2}\right)\right]^{(-1)^{m+n+1}}\times\prod_{m=2q+1}^{2q+n-1}\left[\cos^{2}\left(\frac{pm\pi}{q}+\frac{\phi}{2}\right)\right]^{(-1)^{m+n+1}}$ (15)
$=$ $\prod_{m=1}^{2q}\left[\cos^{2}\left(\frac{pm\pi}{q}+\frac{\phi}{2}\right)\right]^{(-1)^{m+n+1}}\times\prod_{m=1}^{n-1}\left[\cos^{2}\left(\frac{pm\pi}{q}+\frac{\phi}{2}+2p\pi\right)\right]^{(-1)^{m+n+1}},$ (16)
which gives
${\alpha_{n+2q}}=$ $\prod_{m=1}^{2q}\left[\cos^{2}\left(\frac{pm\pi}{q}+\frac{\phi}{2}\right)\right]^{(-1)^{m+n+1}}\times{\alpha_{n}},$ (17)
Hence, to demonstrate the periodicity of ${\alpha_{n}}$, one needs to demonstrate that
$\prod_{m=1}^{2q}\left[\cos^{2}\left(\frac{pm\pi}{q}+\frac{\phi}{2}\right)\right]^{(-1)^{m}}\overset{?}{=}1.$ (18)
or equivalently that
$\left|\prod_{m=1}^{q}\cos\left(\frac{2mp\pi}{q}+\frac{\phi}{2}\right)\right|\overset{?}{=}\left|\prod_{m=1}^{q}\cos\left(\frac{2mp\pi}{q}+\frac{\phi}{2}+\frac{\pi}{q}\right)\right|.$ (19)
Let us consider the Bézout’s identity $ax+by=1$, which has integer solutions if $x,y$ are coprimes. Thus, for $x=2$ and $y=q$, which are coprimes ($q$ is odd by assumption), one gets that $12l+kq=$ 1 has integer solution $l,k$. Moreover, the integer $k$ is odd; otherwise, $l+kq/2$ is an integer, which cannot be since $l+kq/2=1/2$. Hence
$\prod_{m=1}^{q}\cos\left(\frac{2mp\pi}{q}+\frac{\phi}{2}+\frac{\pi}{q}\right)=\prod_{m=1}^{q}\cos\left(\frac{2mp\pi}{q}+\frac{\phi}{2}+\frac{2\pi l}{q}+k\pi\right)=(-1)^{k}\prod_{m=1}^{q}\cos\left(\frac{2(mp+l)\pi}{q}+\frac{\phi}{2}\right).$ (20)
Now, since the order of the factors in the product is irrelevant, one can rearrange them by redefining $m$ and obtain
$\prod_{m=1}^{q}\cos\left(\frac{2mp\pi}{q}+\frac{\phi}{2}+\frac{\pi}{q}\right)=-\prod_{m=1}^{q}\cos\left(\frac{2m\pi}{q}+\frac{\phi}{2}\right),$ (21)
which thus ends the proof.
Notice that this proof cannot be extended to the case where $q$ is even since, in that case, the equation $12l+kq=$ 1 has no integer solutions. Hence, assuming $\omega=2\pi p/q$ with $p,q$ coprimes, $\alpha_{n}$ is periodic in the lattice index $n$ with period $2q$ when $q$ is odd. Conversely, $\alpha_{n}$ is not, in general, periodic when $q$ is even.